### v1

@@ Abstract
In this paper, we present a new ensemble-based filter method by reconstructing the analysis step of the particle filter through a transport map, which directly transports prior particles to posterior particles. The transport map is constructed through an optimization problem described by the Maximum Mean Discrepancy loss function, which matches the expectation information of the approximated posterior and reference posterior. The proposed method inherits the accurate estimation of the posterior distribution from particle filtering. To improve the robustness of Maximum Mean Discrepancy, a variance penalty term is used to guide the optimization. It prioritizes minimizing the discrepancy between the expectations of highly informative statistics for the approximated and reference posteriors. The penalty term significantly enhances the robustness of the proposed method and leads to a better approximation of the posterior. A few numerical examples are presented to illustrate the advantage of the proposed method over ensemble Kalman filter.
keywords:
Ensemble Assimilation method, Transport Map, Maximum Mean Discrepancy
@@ 1 Introduction
Data assimilation aims to obtain a reliable state estimate of the dynamical system by best combining observations and mathematical models [16, 19, 23]. The development of data assimilation provides powerful tools for applications in many fields, such as geophysics, atmosphere science, and numerical weather prediction [1, 10, 14, 33]. The filtering method is an important class of sequence data assimilation technique. In the Bayesian framework, the main procedure of filtering consists of two steps: the prediction step and the analysis step. The prediction step gives prior information based on the numerical simulation of the transition model, and the analysis step solves a Bayesian inverse problem based on the Bayes’ rule conditioned on given observations.
@@
The main challenge of the filtering method lies in the analysis step because it is difficult to accurately estimate the posterior distribution when the dynamical system is characterized by a high dimensional state space [3, 34], strong nonlinearly transition and observation operator [15, 20], and sparse observations [28]. A theoretical analysis of optimal nonlinear filtering is investigated in [17, 18]. The optimal mean square estimate in filtering is equivalent to the posterior mean. However, it is difficult to determine the exact posterior without particular assumptions on the structure of the filter problem. The ensemble Kalman filter (EnKF) is the most popular data assimilation method for nonlinear systems, which gives a sub-optimal state estimation for weakly nonlinear systems based on Gaussian assumptions [6, 7, 8]. However, the EnKF ignores the inconsistency of the sensitivity of each particle under a strongly nonlinear system [27]. Moreover, due to its Gaussian assumption, it can not correctly approximate the non-Gaussian prior distribution of the state, which significantly limits the application of the EnKF [21]. Another well-known filtering method is particle filtering (PF), which is a Monte Carlo method based on sequential importance sampling. In the asymptotic sense, the posterior approximated by PF converges to the true posterior of the data assimilation problem as the number of particles increases to infinity [9, 16]. However, PF suffers from the curse of dimensionality, which often manifests as particle collapse in sequence updating. The number of particles required in PF grows exponentially with respect to the dimension of state space [2]. Resampling can alleviate the challenge by resetting the particle weights to be the same, but it is often not enough to offset this shortcoming. Local particle filtering methods incorporate the concept of localization into particle filtering, which reduces the risk of particle collapse [25, 29]. However, it is challenging to effectively and consistently merge the locally updated particles.
@@
Although Bayesian filtering methods have different forms, they define explicit or implicit transport maps, which push forward the prior distribution to the posterior distribution based on different assumptions on the prior distribution and the likelihood function. The idea of transferring probability measures through a transport map has been well-developed and has been widely used in fluid dynamics, economics, statistics, machine learning, and many other fields [5, 22, 30]. In recent years, the study of transport has gained much attention to treat the challenges of data assimilation in high-dimensional and non-Gaussian nonlinear systems. Some suitable transport maps can been constructed by optimizing the KL divergence between the distribution of the prior and the distribution corresponding to the inverse transport map of the posterior. Spantini et al., [28], proposed an ensemble filtering method based on the coupling technique and obtained a non-Gaussian extension of the ensemble Kalman filter by constructing the Knothe-Rosenblatt rearrangement. Pulido et al., [23], construct a sequence of kernel embedded mappings that represent a gradient flow based on the principles of local optimal transport. Hoang et al., [13], incorporated machine learning into the conditional mean filter and got a nonlinear transformation based on the condition mean’s orthogonal projection property.
@@
In this paper, we propose a transport map that directly transports the prior particles to the posterior particles by reconstructing the PF analysis step. Unlike implicit sampling filtering methods [4, 34], we convert the construction of the transport map into an optimization problem for the Maximum Mean Discrepancy (MMD) loss function by matching the moment information of the approximated posterior generated by transport map and the posterior of PF. In this way, the weight information of the analysis step of PF is converted into updates of the particle states. The proposed method is called as an ensemble transport filter. Here, we take the function space of MMD as a Reproducing Kernel Hilbert Space (RKHS), and then the MMD can be calculated with kernel function in a simplified way. When the kernel function is a universal kernel, it ensures that the distribution obtained through the transport map converges to the reference posterior given by PF as the MMD loss tends to zero. To enhance the robustness of MMD in handling high-dimensional state space [24], we introduce a penalty term based on the variances of the source and target distributions to push the method to prioritize minimizing the mean discrepancy on highly informative statistics. The variance penalty term effectively enhances the robustness of our method in approximating the filtering posterior, making it more effective and stable. Numerical examples show that the ensemble transport filter with a variance penalty term significantly improves robustness in high-dimensional and non-Gaussian problems.
@@
The rest of the article is organized as follows. In Section 2, we introduce the nonlinear filtering problem of a stochastical dynamical system and build a connection with the transfer of distributions. The transport map is constructed in Section 3 by optimizing the mean discrepancy between the approximated posterior and reference posterior. To enhance the robustness of MMD, a variance penalty term is introduced. Numerical examples are presented to evaluate the performance of the ensemble transport filter in Section 4. Finally, we briefly conclude this article in Section 5.
@@ 2.1 State estimation of dynamical system
$\mathrm{d}\boldsymbol{X}_{t}=\mathcal{M}(\boldsymbol{X}_{t})\mathrm{d}t+\mathrm{d}\mathbf{N}_{\boldsymbol{x}},$
where the model noise $\mathbf{N}_{\boldsymbol{x}}$ is a $\mathbb{R}^{n}$-valued stochastic process. Most often, $\mathbf{N}_{\boldsymbol{x}}$ is set to be a Brownian motion, but sometimes Levy processes are used to account for non-Gaussian noise [28]. The system is discretized in time for simulation. Suppose the discrete stepsize of time is $\Delta t$ , we get $\boldsymbol{X}_{(k)\Delta t}=\boldsymbol{X}_{(k-1)\Delta t}+\mathcal{M}(\boldsymbol{X}_{(k-1)\Delta t})\Delta t+\Delta\mathbf{N}_{\boldsymbol{x}}$ with Euler-Maruyama scheme[12]. This leads to a discrete stochastic dynamical model,
(1) $\boldsymbol{X}_{k}=\mathcal{M}_{k}(\boldsymbol{X}_{k-1})+\boldsymbol{\eta}_{k},$
where $\mathcal{M}_{k}$ is the transition model and $\boldsymbol{\eta}_{k}$, with $k=\{0,1,2,\cdots\}$, are the model noise that is assumed to be statistical independent. Noisy observations of state $\boldsymbol{x}_{k}$ are available at discrete times
(2) $\boldsymbol{Y}_{k}=\mathcal{H}_{k}(\boldsymbol{x}_{k})+\boldsymbol{\epsilon}_{k}.$
where $\mathcal{H}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is the observation operator and $\boldsymbol{\epsilon}_{k}$ is the observation noise.
The discrete representation of dynamical model Eq. 1 and observation operator Eq. 2 can be seen as a state-space model. Fig. 1 shows the directed probability graph of state space model. With the assumed random nature of both model and observation noise, the model states and observations can be described as random variables. Filtering aims to infer the conditional posterior $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$ sequentially, and then a Bayesian framework can be used.
[…]
Figure 1: Directed probability graph of dynamical model Eq. 1 and observation operator Eq. 2 in the form of state-space model
@@ 2.2 Sequential Bayesian Filtering
Define the sequence of model states and observations by $\boldsymbol{x}_{0:K}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{K}\}$ and $\boldsymbol{y}_{0:K}=\{\boldsymbol{y}_{1},\boldsymbol{y}_{2},\cdots,\boldsymbol{y}_{K}\}$, respectively. The procedure of sequential Bayesian filtering is to estimate the $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$ iteratively. Assuming the prior of the initial state $\boldsymbol{X}_{0}$ to be $\pi_{\boldsymbol{X}_{0}}$, the sequential Bayesian filtering of the state-space model consists of two steps: forcast and analysis.
@@
    (3) $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}=\int\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{X}_{k-1}}\pi_{\boldsymbol{X}_{k-1}\mid\boldsymbol{Y}_{1:k-1}}\mathrm{d}\boldsymbol{X}_{k-1},$
    where $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{X}_{k-1}}$ is the transition probability.
@@
    (4) $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}=\frac{\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{X}_{k}}}{\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{Y}_{1:k-1}}},$
    where $\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{X}_{k}}$ is the observation likelihood and $\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{Y}_{1:k-1}}$ is the marginal likelihood given by
    $\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{Y}_{1:k-1}}=\int\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{X}_{k}}\mathrm{d}\boldsymbol{X}_{k}.$
@@
The integral of Chapman-Kolmogorov equation Eq. 3 and marginal likelihood $\pi_{\boldsymbol{Y}_{k}\mid\boldsymbol{Y}_{1:k-1}}$ of Eq. 4 can not be calculated analytically in general settings. To overcome this difficulty, an ensemble of particles are used to represent the uncertainty of states and observations in ensemble-based filters. In ensemble-based filters, the filtering process involves transporting particles between three related probability distributions (see Fig. 2).
@@
In the forecast step, the transition density $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{X}_{k-1}}$ is given by a dynamical model, such as like Eq. 1. By an ensemble of particles $\{\boldsymbol{x}_{k-1}^{(i)}\}_{i=1}^{N}$ drawn from posterior distribution $\pi_{\boldsymbol{X}_{k-1}\mid\boldsymbol{Y}_{1:k-1}}$ of the previous time step, we can sample exactly from transition density $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{X}_{k-1}}$ via the discrete dynamical model Eq. 1. Thus, the generated particles $\{\hat{\boldsymbol{x}}_{k}^{(i)}\}_{i=1}^{N}$ can be regarded as samples drawn from prior distribution $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}$ of the current time step. Given the likelihood function and prior distribution approximated by the ensemble of particles, the main task of the analysis step is to seek an approximation of posterior distribution at the current time step. EnKF and its variants approximate the posterior distribution with an ensemble of particles generated from a linear transformation of prior particles under a Gaussian ansatz on prior [27]. Particle filter changes the weights of prior particles through the likelihood function to obtain an accurate estimate of the posterior distribution without altering the state of the prior particles, and a resampling procedure is used to normalize weights [2].
[…]
@@
Figure 2: Propagation of particles between distributions in ensemble-based filtering
@@
For the convenience of presentation, we denote $\hat{\boldsymbol{X}}_{k}$ as the random varible of forecast distribution $\pi_{{\boldsymbol{X}}_{k}\mid\boldsymbol{Y}_{1:k-1}}$. From the point of view of transport, the ensemble-based filter has two transport processes of distribution, which are denoted as push-forward operators $\mathcal{M}_{\sharp}$ and $\mathcal{L}_{\sharp}$ shown in Fig. 2. The $\mathcal{M}_{\sharp}$ and $\mathcal{L}_{\sharp}$ define the transition probability $\pi_{{\boldsymbol{X}}_{k}\mid\boldsymbol{X}_{k-1}}$ and $\pi_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{Y}_{k}}$, respectively,
(5) $\pi_{\hat{\boldsymbol{X}}_{k}}=\int\pi_{\hat{\boldsymbol{X}}_{k}\mid\boldsymbol{X}_{k-1}}\pi_{\boldsymbol{X}_{k-1}\mid\boldsymbol{Y}_{1:k-1}}\mathrm{d}\boldsymbol{X}_{k-1}=\mathcal{M}_{\sharp}\pi_{\boldsymbol{X}_{k-1}\mid\boldsymbol{Y}_{1:k-1}},$
(6) $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}=\int\pi_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{Y}_{k}}\pi_{\hat{\boldsymbol{X}}_{k}}\mathrm{d}\hat{\boldsymbol{X}}_{k}=\mathcal{L}_{\sharp}\pi_{\hat{\boldsymbol{X}}_{k}}.$
@@
Different ensemble filtering methods provide approximate alternatives to the transition probability for the analysis step through different transport maps of ensemble particles. EnKF explicitly constructs a linear transport map, which performs efficiently and robustly in linear and weakly nonlinear systems but performs poorly in strongly nonlinear systems. On the other hand, the particle filter can be viewed as an implicit transport map, providing exact filtering but susceptible to particle degeneracy, thus suffering from the curse of dimensionality. In this article, we intend to obtain an approximation of transition probabilities $\pi_{\boldsymbol{X}_{k}\mid\hat{\boldsymbol{X}}_{k},\boldsymbol{Y}_{k}}$ by constructing an explicit transport map which reconstructs the analysis step of the particle filter.

### v2

@@ Abstract
In this paper, we present a new ensemble-based filter method by reconstructing the analysis step of the particle filter through a transport map, which directly transports prior particles to posterior particles. The transport map is constructed through an optimization problem described by the Maximum Mean Discrepancy loss function, which matches the expectation information of the approximated posterior and reference posterior. The proposed method inherits the accurate estimation of the posterior distribution from particle filtering while gives an extension to high dimensional assimilation problems. To improve the robustness of Maximum Mean Discrepancy, a variance penalty term is used to guide the optimization. It prioritizes minimizing the discrepancy between the expectations of highly informative statistics for the reference posteriors. The penalty term significantly enhances the robustness of the proposed method and leads to a better approximation of the posterior. A few numerical examples are presented to illustrate the advantage of the proposed method over ensemble Kalman filter.
keywords: Ensemble Assimilation method, Transport Map, Maximum Mean Discrepancy.
@@ 1 Introduction
Data assimilation aims to obtain a reliable state estimate of the dynamical system by best combining observations and mathematical models [16, 19, 23]. This approach has evolved into a powerful tool with applications across various fields, including geophysics, atmosphere science, and numerical weather prediction [2, 10, 14, 36]. A key category within sequential data assimilation techniques is the filtering method. In the Bayesian framework, filtering consists of two primary steps: the prediction step and the analysis step. The prediction step provides prior information by simulating the transition model numerically, while the analysis step applies Bayes’ rule to solve a Bayesian inverse problem, incorporating observational data to refine the state estimate.
@@
The main challenge of the filtering method lies in the analysis step since it is difficult to accurately estimate the posterior distribution when the dynamical system is characterized by a high dimensional state space [4, 38], strong nonlinearly transition and observation operator [15, 20], and sparse observations [29]. A theoretical analysis of optimal nonlinear filtering is investigated in [18, 17]. The optimal mean square estimate in filtering is equivalent to the posterior mean. However, it is difficult to determine the exact posterior without particular assumptions on the structure of the filter problem. The ensemble Kalman filter (EnKF) is the most popular data assimilation method for nonlinear systems, which gives a best linear unbias state estimation (BLUE) for weakly nonlinear systems based on a linear transformation [6, 7, 8]. However, the EnKF ignores the inconsistency of the sensitivity of each particle under a strongly nonlinear system [27]. Moreover, it can not correctly approximate the non-Gaussian prior distribution of the state, which significantly limits the application of the EnKF [21]. Another well-known filtering method is particle filtering (PF), which is a Monte-Carlo method based on sequential importance sampling. In the asymptotic sense, the posterior approximated by PF converges to the true posterior of the data assimilation problem as the number of particles increases to infinity [9, 16]. However, PF suffers from the curse of dimensionality, which often manifests as particle collapse in sequence updating. The number of particles required in PF grows exponentially with respect to the dimension of state space [3]. Resampling can alleviate the challenge by resetting the particle weights to be the same, but it is often not enough to offset this shortcoming [28]. In high-dimensional systems, minor differences in states can result in particle weights differing by several orders of magnitude, and resampling may cause the ensemble to collapse into just a few particles. Local particle filtering methods incorporate the concept of localization into particle filtering, which reduces the risk of particle collapse [25, 30]. However, it is challenging to effectively and consistently merge the locally updated particles across domains.
@@
Although Bayesian filtering methods vary in form, they all define explicit or implicit transport maps that propagate the prior distribution to the posterior distribution, based on different assumptions regarding the prior distribution and the likelihood function. The idea of transferring probability measures through a transport map has been well-developed and widely used in fluid dynamics, economics, statistics, machine learning, and many other fields [5, 22, 32]. In recent years, the study of transport has gained much attention to treat the challenges of data assimilation in high-dimensional and non-Gaussian nonlinear systems. Some suitable transport maps can been constructed by optimizing the KL divergence between the distribution of the prior and the distribution corresponding to the inverse transport map of the posterior. Spantini et al., [29], proposed an ensemble filtering method based on the coupling technique and obtained a non-Gaussian extension of the ensemble Kalman filter by constructing the Knothe-Rosenblatt rearrangement. Likelihood-free algorithms are proposed by leveraging a block-triangular structure of transport maps and samples from the joint distribution [1, 34]. Pulido et al., [23], construct a sequence of kernel embedded mappings that represent a gradient flow based on the principles of local optimal transport. Hoang et al., [13], incorporated machine learning into the conditional mean filter and got a nonlinear transformation based on the condition mean’s orthogonal projection property.
@@
In this paper, we propose a transport map that directly transforms the prior particles to the posterior particles by reconstructing the analysis step of PF. We formulate the construction of the transport map as an optimization problem, minimizing the Maximum Mean Discrepancy (MMD) loss function, which matches the moment information of the approximated posterior generated by the transport map with the posterior of the PF. In this way, the weight information of the analysis step of PF is converted into updates for the particle states. The proposed method is called the ensemble transport filter. In this approach, we define the function space of MMD as a Reproducing Kernel Hilbert Space (RKHS), allowing the MMD to be efficiently computed using a kernel function. When the kernel function is a universal kernel, it guarantees that the distribution obtained via the transport map converges to the reference posterior provided by the particle filter as the MMD loss approaches zero. To improve the robustness of MMD in handling high-dimensional state spaces [24], we introduce a penalty term based on the variances of the reference distributions. This encourages the method to prioritize minimizing the mean discrepancy for highly informative statistics. The variance penalty term significantly enhances the method’s robustness in approximating the filtering posterior, making it more effective and stable. Numerical examples demonstrate that the ensemble transport filter with a variance penalty term performs robustly in high-dimensional and non-Gaussian problems. For high-dimensional cases, we locally split the observation into multiple lower-dimensional components. The transport map for the original problem is then constructed by composing the transport maps derived from the separate Bayesian inversions.
@@
The rest of the article is organized as follows. In Section 2, we introduce the nonlinear filtering problem of a stochastic dynamical system and build a connection with the transfer of distributions. The transport map is constructed in Section 3 by optimizing the mean discrepancy between the approximated posterior and reference posterior. To improve the robustness of MMD, a variance penalty term is introduced. Additionally, an extension for high-dimensional problems based on observation splitting is discussed. Numerical examples are presented to evaluate the performance of the ensemble transport filter in Section 4. Finally, we briefly conclude this article in Section 5.
@@ 2.1 State estimation of dynamical system
$\mathrm{d}\bm{X}_{t}=\mathbf{M}(\bm{X}_{t})\mathrm{d}t+\mathrm{d}\mathbf{N}_{\bm{x}},$
where the model noise $\mathbf{N}_{\bm{x}}$ is a $\mathbb{R}^{n}$-valued stochastic process. Most often, $\mathbf{N}_{\bm{x}}$ is set to be a Brownian motion, but sometimes Levy processes are used to account for non-Gaussian noise [29]. The system is discretized in time for simulation. Suppose the discrete stepsize of time is $\Delta t$ , we get $\bm{X}_{(k)\Delta t}=\bm{X}_{(k-1)\Delta t}+\mathbf{M}(\bm{X}_{(k-1)\Delta t})\Delta t+\Delta\mathbf{N}_{\bm{x}}$ with Euler-Maruyama scheme[12]. This leads to a discrete stochastic dynamical model,
$\bm{X}_{k}=\mathbf{M}_{k}(\bm{X}_{k-1})+\bm{\eta}_{k},$ (2.1)
where $\mathbf{M}_{k}$ is the transition model and $\bm{\eta}_{k}$, with $k=\{0,1,2,\cdots\}$, are the model noise that is assumed to be statistical independent. Noisy observations of state $\bm{X}_{k}$ are available at discrete times
$\bm{Y}_{k}=\mathbf{H}_{k}(\bm{X}_{k})+\bm{\epsilon}_{k}.$ (2.2)
where $\mathbf{H}_{k}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is the observation operator and $\bm{\epsilon}_{k}$ is the observation noise.
The discrete representation of dynamical model (2.1) and observation operator (2.2) can be seen as a state-space model. Figure 2.1 shows the directed probability graph of state space model. With the assumed random nature of both model and observation noise, the model states and observations can be described as random variables. Filtering aims to infer the conditional posterior $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ sequentially, and then a Bayesian framework can be used.
Figure 2.1: Directed probability graph of dynamical model (2.1) and observation operator (2.2) in the form of state-space model
@@ 2.2 Sequential Bayesian Filtering
Define the sequence of model states and observations by $\bm{x}_{0:K}=\{\bm{x}_{0},\bm{x}_{1},\cdots,\bm{x}_{K}\}$ and $\bm{y}_{1:K}=\{\bm{y}_{1},\bm{y}_{2},\cdots,\bm{y}_{K}\}$, respectively. The aim of sequential Bayesian filtering is to estimate the $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ iteratively. Assuming the prior of the initial state $\bm{X}_{0}$ to be $\pi_{\bm{X}_{0}}$, the sequential Bayesian filtering of the state-space model consists of two steps: forcast and analysis.
@@
    $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k-1}}=\int\pi_{\bm{X}_{k}\mid\bm{X}_{k-1}}\pi_{\bm{X}_{k-1}\mid\bm{Y}_{1:k-1}}\mathrm{d}\bm{X}_{k-1},$ (2.3)
    where $\pi_{\bm{X}_{k}\mid\bm{X}_{k-1}}$ is the transition probability.
@@
    $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}=\frac{\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k-1}}\pi_{\bm{Y}_{k}\mid\bm{X}_{k}}}{\pi_{\bm{Y}_{k}\mid\bm{Y}_{1:k-1}}},$ (2.4)
    where $\pi_{\bm{Y}_{k}\mid\bm{X}_{k}}$ is the observation likelihood and $\pi_{\bm{Y}_{k}\mid\bm{Y}_{1:k-1}}$ is the marginal likelihood given by
    $\pi_{\bm{Y}_{k}\mid\bm{Y}_{1:k-1}}=\int\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k-1}}\pi_{\bm{Y}_{k}\mid\bm{X}_{k}}\mathrm{d}\bm{X}_{k}.$
@@
The integral of Chapman-Kolmogorov equation (2.3) and marginal likelihood $\pi_{\bm{Y}_{k}\mid\bm{Y}_{1:k-1}}$ of (2.4) are generally intractable to compute analytically under most settings. To overcome this difficulty, ensemble-based filters use an ensemble of particles to represent the uncertainty of states and observations. In ensemble-based filters, the filtering process involves transporting particles between three related probability distributions (see Figure 2.2).
@@
In the forecast step, the transition density $\pi_{\bm{X}_{k}\mid\bm{X}_{k-1}}$ is defined by a dynamical model, such as (2.1). Using an ensemble of particles $\{\bm{x}_{k-1}^{(i)}\}_{i=1}^{N}$ drawn from posterior distribution $\pi_{\bm{X}_{k-1}\mid\bm{Y}_{1:k-1}}$ of the previous assimilation window, samples can be directly drawn from transition density $\pi_{\bm{X}_{k}\mid\bm{X}_{k-1}}$ using the discrete dynamical model (2.1). The resulting particles $\{\hat{\bm{x}}_{k}^{(i)}\}_{i=1}^{N}$ represent samples from prior distribution $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k-1}}$ of the current assimilation window. In the analysis step, the primary goal is to approximate the posterior distribution at the current assimilation window using the likelihood function and the prior distribution, which is represented by the ensemble of particles. EnKF and its variants approximate the posterior distribution with an ensemble of particles generated from a linear transformation of prior particles under a Gaussian ansatz [27, 29]. Particle filter adjusts the weights of prior particles through the likelihood function to obtain an accurate estimate of the posterior distribution without altering the state of the prior particles, and a resampling procedure is used to normalize weights [3].
@@
Figure 2.2: Propagation of particles between distributions in ensemble-based filtering
@@
For clarity, we denote $\hat{\bm{X}}_{k}$ as the random variable of forecast distribution $\pi_{{\bm{X}}_{k}\mid\bm{Y}_{1:k-1}}$. From the perspective of transport, the ensemble-based filter involves two distribution transport processes, represented as push-forward operators $\mathcal{M}_{\sharp}$ and $\mathcal{L}_{\sharp}$, as illustrated in Figure 2.2. The $\mathcal{M}_{\sharp}$ and $\mathcal{L}_{\sharp}$ define the transition probability $\pi_{{\bm{X}}_{k}\mid\bm{X}_{k-1}}$ and $\pi_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{Y}_{k}}$, respectively,
$\pi_{\hat{\bm{X}}_{k}}=\int\pi_{\hat{\bm{X}}_{k}\mid\bm{X}_{k-1}}\pi_{\bm{X}_{k-1}\mid\bm{Y}_{1:k-1}}\mathrm{d}\bm{X}_{k-1}=\mathcal{M}_{\sharp}\pi_{\bm{X}_{k-1}\mid\bm{Y}_{1:k-1}},$ (2.5)
$\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}=\int\pi_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{Y}_{k}}\pi_{\hat{\bm{X}}_{k}}\mathrm{d}\hat{\bm{X}}_{k}=\mathcal{L}_{\sharp}\pi_{\hat{\bm{X}}_{k}}.$ (2.6)
@@
Different ensemble filtering methods provide approximate alternatives to the transition probability for the analysis step through different transport maps of ensemble particles. EnKF explicitly constructs a linear transport map, which is efficient and robust in linear and weakly nonlinear systems but performs poorly in strongly nonlinear systems. On the other hand, the particle filter can be viewed as an implicit transport map, providing exact filtering but susceptible to particle degeneracy, thus suffering from the curse of dimensionality. In this article, we aim to obtain an approximation of transition probabilities $\pi_{\bm{X}_{k}\mid\hat{\bm{X}}_{k},\bm{Y}_{k}}$ by constructing an explicit transport map that reconstructs the analysis step of the particle filter.